. List and discuss, in your own words, the four key parameters of the Michaelis–Menten Equation. Your response must clearly outline the definition of each of the parameters,and its application to enzyme activity/substrate interaction affinity/catalytic efficiency


1
Expert's answer
2021-10-15T01:46:01-0400

Michaelis–Menten Equation is a model used in enzymes kinetics.

The equation is:"𝑉 =\frac{C(Vmax)}{C + Km}"

The four parameters are:

V- The dependent variable.

C- The Independent variable

Vmax is the maximum velocity and may serve as horizontal asymptote.

Km is the ED50 or Michaelis constant .

The Km and Vmax are the two parameters to be calculated.

The rate of enzyme reaction is equal to "\frac{Vmax \times\ concentration}{Km + concentration}"

Therefore, the higher the Km the lower the enzyme reaction as per the above model.